# MOSAIC: Generating Consistent, Privacy-Preserving Scenes from Multiple Depth Views in Multi-Room Environments

### v2

@@ Abstract
We introduce a novel diffusion-based approach for generating privacy-preserving digital twins of multi-room indoor environments from depth images only. Central to our approach is a novel Multi-view Overlapped Scene Alignment with Implicit Consistency (MOSAIC) model that explicitly considers cross-view dependencies within the same scene in the probabilistic sense. MOSAIC operates through a novel inference-time optimization that avoids error accumulation common in sequential or single-room constraint in panorama-based approaches. MOSAIC scales to complex scenes with zero extra training and provably reduces the variance during denoising processes when more overlapping views are added, leading to improved generation quality. Experiments show that MOSAIC outperforms state-of-the-art baselines on image fidelity metrics in reconstructing complex multi-room environments. Project page is available at: https://mosaic-cmubig.github.io.
@@ 1 Introduction
Autonomous scene reconstruction is a crucial capability in robotics and computer vision [22, 23, 15, 1], with applications spanning virtual reality, architectural design, and AI-driven simulation. However, existing multi-view 3D reconstruction methods rely heavily on capturing RGB images, which poses significant privacy risks in sensitive environments such as hospitals, factories, elder care facilities, and private residences. Capturing RGB data in these settings can inadvertently expose personal information [29, 44], limiting the practicality of generative AI solutions for real-world deployments where visual privacy must be preserved.
@@
A promising solution for privacy-sensitive environments involves deploying mobile robots that collect only geometric data, such as depth images: this strategy preserves the scene’s structural layout while preventing the capture of sensitive texture information. In order to create digital replicas of these scenes, we need algorithms that are capable of generating high quality, multiview-consistent sequential images that align with ground truth geometry, while maintaining visual coherence across viewpoints.
@@
Various approaches for scene level multiview image generation have been proposed, with autoregressive generation techniques [13, 8, 42, 41, 7] representing the current mainstream trend. These approaches leverage powerful inpainting models to iteratively render unseen parts of the scene from sequential camera perspectives. However, these approaches exhibit style drift, where images generated for initial viewpoints often differ stylistically from those produced when revisiting the same location from different angles, creating visual inconsistencies. Moreover, warp-and-inpainting methods frequently encounter depth misalignment issues, leading to compounding errors over sequential generations [33, 21]. Unlike prior methods that perform diffusion synchronization tasks [2, 18, 9, 20, 16] in constrained settings (limited viewpoint variations), our task requires coherence between viewpoints based on arbitrary trajectory features, extensive perspective changes, and spatial discontinuities. This fundamental distinction makes existing methods ineffective for complex multi-room environments with widely-varying camera positions.
To address these limitations, we introduce Multi-view Overlapped Scene Alignment with Implicit Consistency (MOSAIC), a novel approach that achieves multi-view consistent RGB image generation from depth data, without compromising privacy in scene level, and while remaining robust to arbitrary egocentric perspective changes across multi-room environments. Rather than requiring extensive training on arbitrary trajectories, MOSAIC operates through a novel multiview inference-time optimization that ensures cross-view consistency during the diffusion process. By minimizing projection differences between overlapping views at each denoising step, our method guarantees visual coherence across the entire scene. A key advantage of MOSAIC is its ability to reduce variance and improve image-depth alignment as more overlapping views are added—the opposite behavior to current autoregressive approaches, which accumulate errors over time. Moreover, we propose a depth-weighted optimization, which naturally prioritizes information from optimal viewpoints, and a pixel-space refinement mechanism which ensures that latent-space consistency translates to visual harmony in the final output. Importantly, MOSAIC scales effortlessly to complex multi-room environments without additional training, making it practical for complex real-world applications. Our experiments demonstrate superior performance over existing methods across multiple metrics, establishing a new state-of-the-art for scene level multi-view image generation with geometric priors.
Our contributions are summarized as follows:
  1. 1.
    We propose MOSAIC, a novel diffusion-based model for multi-view-consistent image generation via explicit modeling of cross-view dependencies.
  2. 2.
    We propose a novel multi-view inference-time optimization as a practical sampling strategy that scales to complex scenes and more depth views with no extra training.
  3. 3.
    We show theoretically that with more overlapping depth views, MOSAIC reduces the variance of generation process, mitigating error accumulations in current autoregressive multi-view image generation pipelines.
  4. 4.
    We introduce depth-weighted projection loss and pixel-space refinement to enforce visual coherence. Extensive experiments demonstrate that MOSAIC generalizes across diverse viewpoints, significantly outperforming state-of-the-art methods in generation quality.
@@ 2 Related Work
Diffusion-Guided Scene-Level Multi-view Generation. Recent advances in 2D diffusion-based generative models [32, 11, 30, 43] have catalyzed significant progress in scene-level multiview image generation—particularly in the use of autoregressive generation techniques that leverage warping and inpainting mechanisms for, e.g., scene completion [13, 8, 42, 41, 7]. These mechanisms can introduce style drift errors where images in later iterations differ stylistically from the initial ones, leading to artifacts and distortions. Alternative approaches attempt to maintain style consistency by generating panoramic representations using off-the-shelf diffusion models [33, 19, 31, 37]. While these methods achieve internal coherence, they cannot handle multiple rooms with complex layouts or flexible camera trajectories, limiting their practicality in the real world. Recent video diffusion models demonstrate promising capabilities for generating temporally consistent image sequences  [4, 12, 40, 45]. However, they do not apply to our task due to the inability to incorporate depth conditions and limited tolerance against perspective changes. Our method, on the other hand, explicitly models cross-view consistency and generalizes to arbitrary viewpoints without imposing layout constraints.
@@
Diffusion Synchronization. Several works have explored synchronization techniques for diffusion models [2, 18, 9, 20, 16] to achieve consistency across multiple outputs. However, existing approaches primarily address specialized cases: panoramic view generation [2, 18, 16], orthogonal transformations [9], or object-level consistency with evenly distributed camera positions and well established UV map [16, 20]. Our work addresses a fundamentally different challenge: maintaining consistent image generation across scene-level camera trajectories with arbitrary viewpoints and significant perspective variations. In this complex setting, naive averaging of denoising features fails to preserve cross-view consistency due to the non-orthogonal nature of the transformations and the extensive variance in viewpoint geometry.
@@
Diffusion Inference-Time Optimization. To circumvent the need for large-scale model fine-tuning, several approaches have explored test-time optimization with directional guidance. Diffusion-TTA [26] adapts discriminative models using pre-trained generative diffusion models, updating parameters through gradient backpropagation at inference time. Other methods [24, 35, 25] compute directional controls for single-instance generation (music, images) by calculating directional losses during diffusion and back-propagating to update denoising features. While these methods focus on intra-level control with easily quantifiable directions, our approach is the first to propose multi-channel test-time optimization for cross-view control using self-provided directional guidance. This novel approach enables us to maintain consistency across arbitrary viewpoints without requiring additional training, addressing the fundamental limitations of existing methods when applied to complex multi-room environments.
Figure 2: MOSAIC Overview. MOSAIC implements a multi-channel denoising framework, as illustrated in (a). We formulate the training objective and address it through projection loss with inference-time optimization (b). In the final denoising phase, a pixel space refinement loss enforces precise RGB consistency constraints (c).
@@ 3 Problem Definition
We seek to acquire digital replicas of real-world, multi-room indoor environments in a privacy-preserving manner. To safeguard privacy, we refrain from collecting sensitive real RGB data and capture only geometric structures instead. From these structures, we generate synthetic RGB data—producing complex, photorealistic scenes that closely align with reality while ensuring no actual environmental details are directly revealed. In this work, we assume that the geometric structures are collected as multiple depth images $\{d^{[1]},\dots,d^{[N]}\}$ by mobile robots deployed to real-world scenes. To fully cover multi-room environments, the mobile robot plans proper camera pose sequences to ensure overlapping views. Namely, for each $d^{[i]}$, there must exist some area that is also covered by at least one other view $d^{[j]}$ $(j\neq i)$. From $\{d^{[i]}\}_{i\in[N]}$, we are interested in generating corresponding RGB images $\{x^{[i]}\}_{i\in[N]}$ that are multi-view consistent: overlapping depth views must lead to consistent RGB outputs where they overlap in ${\bar{x}}$. Namely, for each pixel $p$ in ${\bar{x}}$ that is covered by multiple depth views, i.e., $I_{p}:=\{i\in[N]\mid p\in d^{[i]}\}$, the RGB output should agree:
@@ 4.1 Preliminaries
Denoising Diffusion Implicit Models (DDIM) [32] train a generative model $p_{\theta}(z_{0})$ to approximate a data distribution $q(z_{0})$ given samples $z\in\mathcal{Z}$ from $q$. DDIM considers the following non-Markovian inference model:
@@
where $\sigma\in\mathbb{R}_{\geq 0}^{T}$ is a real vector and $z_{1:T}$ are latent variables in $\mathcal{Z}$. The inference process $q_{\sigma}\left(z_{t-1}\mid z_{t},z_{0}\right)$ is parameterized by a decreasing sequence $\alpha_{1:T}\in(0,1]^{\top}$. To approximate $q(z_{0})$, DDIM learns a generative process
@@
In this work, we generate RGB images from depth views $d$, via ControlNet [43]: $\epsilon_{\theta}^{(t)}(z_{t},d)$. In addition, we sample latent RGB representations $z$ from $p_{\theta}$ which can later be decoded into full images via a pre-trained VAE decoder [17].
@@
For clarity, we drop the depth dependency $d^{[i]}$ in notations in the rest of this paper. While the generation of a single image $x_{i}$ can be readily solved via depth-conditioned diffusion models [43], the generation of complete scenes from multiple depth images remains unsolved.
@@ 4.2 MOSAIC Formulation
Intuitively, the sampling procedure in Eq. 6 would generate inconsistent RGB views because the generative process $p_{\theta}$ is trained with independent samples $z_{0}\sim q(z_{0})$. In our case, however, the samples $z_{0}^{[1:N]}$ are indeed dependent since they are taken from the same complete scene ${\bar{z}_{0}}$. To fundamentally address the consistency issue, our key insight is to explicitly model such dependency by incorporating extra projection-based conditionals $q(z_{0}^{[i]}\mid{\bar{z}_{0}})$ in the inference procedures. In practice, given ${\bar{z}_{0}}$, $z_{0}^{[i]}$ can be collected by simply sampling a camera pose with projection function $\pi^{[i]}$ and setting $z_{0}^{[i]}=\pi^{[i]}({\bar{z}_{0}})$; subsequent latents $z_{1:T}^{[i]}$ can be sampled as usual. The updated inference procedure for $N$ depth views from the same scene ${\bar{z}_{0}}$ can be rewritten from Eq. 2 as:
@@ 4.3 Sampling from MOSAIC with Multi-Channel Inference-Time Optimization
Although it is impractical to directly train $p_{\theta,\phi}$ using Eq. 13, it is possible to approximate the samples from $p_{\theta,\phi}$ by fine-tuning the output of pre-trained models $p_{\theta}$, e.g., Latent Diffusion Model (LDM) [30]. We observe that for MOSAIC, $p_{\theta,\phi}$ fits trajectories $z_{1:T}^{[1]},\dots,z_{1:T}^{[N]}$ that are mutually dependent at each step $t=1,\dots,T$ through $z_{0}^{1:N}$, while such dependency is missing from LDM $p_{\theta}$. Furthermore, this dependency is exactly captured by $J_{\mathrm{Proj}}(\phi)$. If $p_{\theta,\phi}^{*}$ optimally solves $J_{\sigma,\mathrm{MOSAIC}}$ and $z_{0}^{[1:N]}\sim p_{\theta,\phi}^{*}$ , there must exist some ${\bar{z}_{0}}$ and corresponding projections $\pi^{[1:N]}$ such that $z_{0}^{[i]}=\pi^{[i]}({\bar{z}_{0}}),~{}\forall i\in[N]$. Then, $J_{\mathrm{Proj}}(\phi)$ must have been minimized to 00 with $f_{\phi^{*}}^{[i]}=(\pi^{[i]})^{-1},~{}\forall i$. In other words, samples from an ideal MOSAIC model should have zero projected cross-view error in expectation as long as $f_{\phi}^{[i]}$ exactly matches the inverse projection $(\pi^{[i]})^{-1}$. While it is generally hard to fit an inverse projection, it is unnecessary in our case because the ground truth projections $\pi^{[1:N]}$ are directly available when collecting depth conditions $d^{[1:N]}$ [^1] With mobile robots, $\pi^{[i]}$ can be computed from the camera/robot poses in the world coordinate when taking each depth view $d^{[i]}$..
Hence, we can approximately sample from MOSAIC by fine-tuning LDM (i.e., $p_{\theta}$) output at each denoising step $t-1$ by solving the following inference-time optimization through gradient decent:
@@
$L_{\text{Proj}}^{\text{DW}}(z_{t}^{[1:N]})=\sum_{i,j}w_{i,j}\cdot\|(\pi^{[i]})^{-1}(\hat{z}_{0}^{[i]})-(\pi^{[j]})^{-1}(\hat{z}_{0}^{[j]})\|$ (16)
where $w_{i,j}$ weights the importance of consistency between views $i$ and $j$ based on their relative depths. To prioritize views with smaller depth values, we employ weighting scheme:

### v3

@@ Abstract
We introduce a diffusion-based approach for generating privacy-preserving digital twins of multi-room indoor environments from depth images only. Central to our approach is a novel Multi-view Overlapped Scene Alignment with Implicit Consistency (MOSAIC) model that explicitly considers cross-view dependencies within the same scene in the probabilistic sense. MOSAIC operates through a multi-channel inference-time optimization that avoids error accumulation common in sequential or single-room constraints in panorama-based approaches. MOSAIC scales to complex scenes with zero extra training and provably reduces the variance during denoising process when more overlapping views are added, leading to improved generation quality. Experiments show that MOSAIC outperforms state-of-the-art baselines on image fidelity metrics in reconstructing complex multi-room environments. Resources and code are at https://mosaic-cmubig.github.io.
@@ 1 Introduction
Autonomous scene reconstruction is a crucial capability in robotics and computer vision [mildenhall2020nerfrepresentingscenesneural, murez2020atlas, kerbl20233dgaussiansplattingrealtime, amatare2024dtradardigitaltwinassisted], with applications spanning virtual reality, architectural design, and AI-driven simulation. However, existing multi-view 3D reconstruction methods rely heavily on capturing RGB images, which poses significant privacy risks in sensitive environments such as hospitals, factories, elder care facilities, and private residences. Capturing RGB data in these settings can inadvertently expose personal information [ravi2021reviewvisualprivacypreservation, zhao2022privacypreservingreflectionrenderingaugmented], limiting the practicality of generative AI solutions for real-world deployments where visual privacy must be preserved.
@@
A promising solution for privacy-sensitive environments involves deploying mobile robots that collect only geometric data, such as depth images: this strategy preserves the scene’s structural layout while preventing the capture of sensitive texture information. In order to create digital replicas of these scenes, we need algorithms that are capable of generating high-quality, multiview-consistent sequential images that align with ground truth geometry, while maintaining visual coherence across viewpoints.
@@
Various approaches for scene-level multi-view image generation have been proposed, with autoregressive generation techniques [hollein2023text2room, fridman2023scenescape, yu2024wonderjourneygoing, yu2024wonderworldinteractive3dscene, chung2023luciddreamer] representing the current mainstream trend. These approaches leverage powerful inpainting models to iteratively render unseen parts of the scene from sequential camera perspectives. However, these approaches exhibit style drift, where images generated for initial viewpoints often differ stylistically from those produced when revisiting the same location from different angles, creating visual inconsistencies. Moreover, warp-and-inpainting methods frequently encounter depth misalignment issues, leading to compounding errors over sequential generations [song2023roomdreamertextdriven3dindoor, controlnetinpaint2023]. Unlike prior methods that perform diffusion synchronization tasks [bartal2023multidiffusionfusingdiffusionpaths, lee2023syncdiffusioncoherentmontagesynchronized, geng2024visualanagramsgeneratingmultiview, liu2023textguidedtexturingsynchronizedmultiview, kim2024synctweediesgeneralgenerativeframework] in constrained settings (limited viewpoint variations), our task requires coherence between viewpoints based on arbitrary trajectory features, extensive perspective changes, and spatial discontinuities. This fundamental distinction makes existing methods ineffective for complex multi-room environments with widely-varying camera positions.
To overcome these limitations, we present MOSAIC (Multi‑view Overlapped Scene Alignment with Implicit Consistency), a training‑free diffusion framework that converts privacy‑preserving depth sequences into photorealistic, geometry‑aligned RGB views across large, cluttered, multi‑room environments. Previous work [tang2023mvdiffusion] enforces consistency by fine‑tuning a modified architecture on trajectory‑specific data, limiting scalability and generalization; MOSAIC instead performs a lightweight multi‑view inference‑time optimization that scales to unseen environments and enforces cross‑view agreement at every denoising step, while preserving the intrinsic knowledge of a large pre‑trained model [podell2023sdxlimprovinglatentdiffusion]. A depth‑weighted projection loss highlights the most reliable viewpoints, and a final pixel‑space refinement converts latent‑space coherence into precisely aligned images. As more overlapping views are added, MOSAIC reduces variance and tightens image–depth alignment, unlike autoregressive pipelines that compound errors, thus delivering stable reconstructions over arbitrarily long trajectories. Our experiments demonstrate superior performance over existing methods across multiple metrics, establishing a new state-of-the-art for scene level multi-view image generation with geometric priors.
In summary, (1) we introduce MOSAIC, a training-free diffusion pipeline that turns arbitrary depth sequences into photorealistic geometry-aligned RGB views, (2) we devise an inference-time sampler, with depth-weighted projection loss and late pixel refinement, that enforces strict cross-view coherence, (3) we theoretically prove that adding overlapping views monotonically reduces denoising variance, curbing error accumulation, and (4) we demonstrate state-of-the-art results in fidelity, prompt-following, and geometry alignment in cluttered, multi-room environments.
@@ 2 Related Work
Diffusion‑guided scene‑level multi‑view generation. Recent advances in 2D diffusion-based generative models [ddim, ddpm, rombach2022high, control_net, podell2023sdxlimprovinglatentdiffusion, esser2024scalingrectifiedflowtransformers] have catalyzed significant progress in scene-level multiview image generation, yet most pipelines rely on iterative warping and inpainting for view completion [hollein2023text2room, fridman2023scenescape, yu2024wonderjourneygoing, yu2024wonderworldinteractive3dscene, chung2023luciddreamer], where style drift accumulates across iterations and produces artifacts. A panorama‑first strategy mitigates drift by re‑projecting a single generated panorama [song2023roomdreamertextdriven3dindoor, li2024scenedreamer360textdriven3dconsistentscene, schult2023controlroom3droomgenerationusing, wang2024roomtextexturingcompositionalindoor], but cannot cope with multi‑room layouts or long, unconstrained trajectories. MVDiffusion [tang2023mvdiffusion] requires training an additional module for cross-view consistency, which limits the generalization to unseen environments and prompt following fidelity. SceneTex [chen2023scenetexhighqualitytexturesynthesis] assumes a pre‑meshed scene and performs score‑distillation texture optimization, requiring $\sim$20 hours per scene and often seeking to a single appearance mode; our training‑free DDIM process finishes in few minutes while preserving generation diversity under single text prompt. Recent video‑diffusion models provide temporal coherence [svd, ho2022video, yang2024cogvideoxtexttovideodiffusionmodels, zheng2024open], but lack depth conditioning and break under large viewpoint changes. Our method, on the other hand, explicitly models cross-view consistency by optimizing MOSAIC objective and generalizes to arbitrary viewpoints without imposing layout constraints.
@@
Diffusion Synchronization. Several works have explored synchronization techniques for diffusion models [bartal2023multidiffusionfusingdiffusionpaths, lee2023syncdiffusioncoherentmontagesynchronized, geng2024visualanagramsgeneratingmultiview, liu2023textguidedtexturingsynchronizedmultiview, kim2024synctweediesgeneralgenerativeframework] to achieve consistency across multiple outputs. However, existing approaches primarily address specialized cases: panoramic view generation [bartal2023multidiffusionfusingdiffusionpaths, lee2023syncdiffusioncoherentmontagesynchronized, kim2024synctweediesgeneralgenerativeframework], orthogonal transformations [geng2024visualanagramsgeneratingmultiview], or object-level consistency with evenly distributed camera positions and well established UV map [kim2024synctweediesgeneralgenerativeframework, liu2023textguidedtexturingsynchronizedmultiview]. Our work addresses a fundamentally different challenge: maintaining consistent image generation across scene-level camera trajectories with arbitrary viewpoints and significant perspective variations.
@@
Diffusion Inference-Time Optimization. To circumvent the need for large-scale model fine-tuning, several approaches have explored test-time optimization with directional guidance. Diffusion-TTA [prabhudesai2023Diffusion-TTA] adapts discriminative models using pre-trained generative diffusion models, updating parameters through gradient backpropagation at inference time. Other methods [novack2024DITTO, tsai2024gdageneralizeddiffusionrobust, phung2023groundedtexttoimagesynthesisattention] compute directional controls for single-instance generation (music, images) by calculating directional losses during diffusion and back-propagating to update denoising features. While these methods focus on intra-level control with easily quantifiable directions, our approach is the first to propose multi-channel test-time optimization for cross-view control using self-provided directional guidance, maintaining consistency across arbitrary viewpoints without requiring additional training.
Figure 2: MOSAIC overview. (a) Multi-channel denoising. Each depth–text–conditioned view is assigned its own latent channel. A shared denoiser iteratively refines the latent set while a multi-channel inference-time optimizer keeps the channels synchronized. (b) Projection loss. At every step the predicted clean latents $z_{0}$ guided depth-weighted projection loss $L_{\text{proj}}$ drives the channels toward a geometry-consistent solution. (c) Pixel-space refinement. During the final denoising stages, the pixel-level loss $L_{\text{pixel}}$ fuses the views and enforces RGB consistency, yielding photorealistic, cross-view-aligned images that can be reconstructed into a coherent 3-D scene.
Figure 3: Qualitative comparison with multi-view baselines. For three indoor scenes (each conditioned on a style prompt) we show: input depth maps, ground-truth RGB, our MOSAIC result, and two baselines (MVDiffusion, Warp-and-Inpainting) sharing the identity input format. Below, we compare against baselines (SceneTex, SceneCraft, Text2Room) using their native inputs. MOSAIC maintains photorealism, cross-view consistency, and prompt fidelity, whereas competing methods exhibit blur, style drift, or geometric artifacts.
@@ 3 Problem Definition
We seek to acquire digital replicas of real-world, multi-room indoor environments in a privacy-preserving manner. To safeguard privacy, we refrain from collecting sensitive real RGB data and capture only geometric structures instead. From these structures, we generate synthetic RGB data—producing complex, photorealistic scenes that closely align with reality while ensuring no actual environmental details are directly revealed. In this work, we assume that the geometric structures are collected as multiple depth images $\{d^{[1]},\dots,d^{[N]}\}$ by mobile robots deployed to real-world scenes. To fully cover multi-room environments, the mobile robot plans proper camera pose sequences to ensure overlapping views. Namely, for each $d^{[i]}$, there must exist some area that is also covered by at least one other view $d^{[j]}$ $(j\neq i)$. From $\{d^{[i]}\}_{i\in[N]}$, we are interested in generating corresponding RGB images $\{x^{[i]}\}_{i\in[N]}$ that are multi-view consistent: overlapping depth views must lead to consistent RGB outputs where they overlap in the 3D scene $\bar{x}$}. Namely, for each pixel $p$ in $\bar{x}$} that is covered by multiple depth views, i.e., $I_{p}:=\{i\in[N]\mid p\in d^{[i]}\}$, the RGB output should agree:
@@ 4.1 Preliminaries
Denoising Diffusion Implicit Models (DDIM) [ddim] train a generative model $p_{\theta}(z_{0})$ to approximate a data distribution $q(z_{0})$ given samples $z\in\mathcal{Z}$ from $q$. DDIM considers the following non-Markovian inference model:
@@
where $\sigma\in\mathbb{R}_{\geq 0}^{T}$ is a real vector and $z_{1:T}$ are latent variables in $\mathcal{Z}$. The inference procedure $q_{\sigma}\left(z_{t-1}\mid z_{t},z_{0}\right)$ is parameterized by a decreasing sequence $\alpha_{1:T}\in(0,1]^{\top}$. To approximate $q(z_{0})$, DDIM learns a generative process
@@
In this work, we generate RGB images from depth views $d$, via ControlNet [control_net]: $\epsilon_{\theta}^{(t)}(z_{t},d)$. In addition, we sample latent RGB representations $z$ from $p_{\theta}$ which can later be decoded into full images via a pre-trained VAE decoder [kingma2022autoencodingvariationalbayes].
@@
For clarity, we drop the depth dependency $d^{[i]}$ in notations in the rest of this paper. While the generation of a single image $x_{i}$ can be readily solved via depth-conditioned diffusion models [control_net], the generation of complete scenes from multiple depth images remains unsolved.
@@ 4.2 MOSAIC Formulation
Intuitively, the sampling procedure in Eq. 6 would generate inconsistent RGB views because the generative process $p_{\theta}$ is trained with independent samples $z_{0}\sim q(z_{0})$. In our case, however, the samples $z_{0}^{[1:N]}$ are indeed dependent since they are taken from the same complete scene $\bar{z}_{0}$}. To fundamentally address the consistency issue, our key insight is to explicitly model such dependency by incorporating extra projection-based conditionals $q(z_{0}^{[i]}\mid{\bar{z}_{0}})$ in the inference procedures. In practice, given the scene in the latent representation $\bar{z}_{0}$}, $z_{0}^{[i]}$ can be collected by sampling a camera pose with projection function $\pi^{[i]}$ and setting $z_{0}^{[i]}=\pi^{[i]}({\bar{z}_{0}})$; subsequent latents $z_{1:T}^{[i]}$ can be sampled as usual. The updated inference procedure for $N$ depth views from the same scene $\bar{z}_{0}$} can be rewritten from Eq. 2 as:
@@ 4.3 Sampling from MOSAIC with Multi-Channel Inference-Time Optimization
Although it is impractical to directly train $p_{\theta,\phi}$ using Eq. 13, it is possible to approximate the samples from $p_{\theta,\phi}$ by fine-tuning the output of pre-trained models $p_{\theta}$, e.g., Latent Diffusion Model (LDM) [rombach2022high]. We observe that for MOSAIC, $p_{\theta,\phi}$ fits trajectories $z_{1:T}^{[1]},\dots,z_{1:T}^{[N]}$ that are mutually dependent at each step $t=1,\dots,T$ through $z_{0}^{1:N}$, while such dependency is missing from LDM $p_{\theta}$. Furthermore, this dependency is exactly captured by $J_{\mathrm{Proj}}(\phi)$. If $p_{\theta,\phi}^{*}$ optimally solves $J_{\sigma,\mathrm{MOSAIC}}$ and $z_{0}^{[1:N]}\sim p_{\theta,\phi}^{*}$ , there must exist some $\bar{z}_{0}$} and corresponding projections $\pi^{[1:N]}$ such that $z_{0}^{[i]}=\pi^{[i]}({\bar{z}_{0}}),~\forall i\in[N]$. Then, $J_{\mathrm{Proj}}(\phi)$ must have been minimized to 0 with $f_{\phi^{*}}^{[i]}=(\pi^{[i]})^{-1},~\forall i$. In other words, samples from an ideal MOSAIC model should have zero projected cross-view error in expectation as long as $f_{\phi}^{[i]}$ exactly matches the inverse projection $(\pi^{[i]})^{-1}$. While it is generally hard to fit an inverse projection, it is unnecessary in our case because the ground truth projections $\pi^{[1:N]}$ are directly available when collecting depth conditions $d^{[1:N]}$ [^1] With mobile robots, $\pi^{[i]}$ can be computed from the camera/robot poses in the world coordinate when taking each depth view $d^{[i]}$..
Hence, we can approximately sample from MOSAIC by fine-tuning LDM (i.e., $p_{\theta}$) output at each denoising step t1t-1 by solving the following inference-time optimization through gradient decent:
@@
$L_{\text{Proj}}^{\text{depth}}(z_{t}^{[1:N]})=\sum_{i,j}w_{i,j}\cdot\|(\pi^{[i]})^{-1}(\hat{z}_{0}^{[i]})-(\pi^{[j]})^{-1}(\hat{z}_{0}^{[j]})\|$ (16)
where $w_{i,j}$ weights the importance of consistency between views $i$ and $j$ based on their relative depth values, weights are calculated per pixel. To prioritize views with smaller depth values, we employ weighting scheme: